1. List all elements in A. A={xϵN:(3x2-x)<55}


1
Expert's answer
2022-02-24T06:23:01-0500

Let us list all elements in "A=\{x\in\N:(3x^2-x)<55\}."

For this let us solve the inequality "3x^2-x<55."

The discriminant of the equation "3x^2-x-55=0" is equal to "D=1+3\cdot4\cdot 55=661," and hence the equation has the solutions "x_1=\frac{1-\sqrt{661}}{6}\approx -4.11" and "x_2=\frac{1+\sqrt{661}}{6}\approx 4.45"

It follows that the real solutions of the inequlity "3x^2-x<55" belong to the set "(\frac{1-\sqrt{661}}{6},\frac{1+\sqrt{661}}{6})."

This set contains the following natural numbers: "1,2,3,4."

Consequently, "A=\{1,2,3,4\}."